OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1

OpenStudy (anonymous):

could it be: \[y+(69/10)x-51/10=0\]

OpenStudy (anonymous):

CHOICES A. 69x + 10y = 51 B. -69x + 10y = 51 C. 69x + 10y = -51 D. -69x + 10y = -51